Entries for Factory Costs and Jobs Completed

Collegiate Publishing Inc. began printing operations on March 1. Jobs 301 and 302 were completed during the month, and all costs applicable to them were recorded on the related cost sheets. Jobs 303 and 304 are still in process at the end of the month, and all applicable costs except factory overhead have been recorded on the related cost sheets. In addition to the materials and labor charged directly to the jobs, $4,500 of indirect materials and $8,200 of indirect labor were used during the month. The cost sheets for the four jobs entering production during the month are as follows, in summary form:

Job 301 Job 302
Direct materials $12,500 Direct materials $18,750
Direct labor 31,000 Direct labor 42,200
Factory overhead 7,750 Factory overhead 10,550
Total $51,250 Total $71,500
Job 303 Job 304
Direct materials $9,940 Direct materials $14,310
Direct labor 16,500 Direct labor 17,100
Factory overhead Factory overhead

Journalize the summary entry to record each of the following operations for March (one entry for each operation):

a. Direct and indirect materials used. If an amount box does not require an entry, leave it blank.

Work in Process
Factory Overhead   
Materials

Feedback

b. Direct and indirect labor used. If an amount box does not require an entry, leave it blank.

     
  

Feedback

c. Factory overhead applied to all four jobs (a single overhead rate is used based on direct labor cost). If an amount box does not require an entry, leave it blank.

Feedback

d. Completion of Jobs 301 and 302. If an amount box does not require an entry, leave it blank.
Answer #1
Overhead rate 25% =7750/31000
a Work in Process 55500 =12500+18750+9940+14310
Factory Overhead 4500
      Materials 60000
b Work in Process 106800 =31000+42200+16500+17100
Factory Overhead 8200
     Wages Payable 115000
c Work in Process 26700 =106800*25%
      Factory Overhead 26700
d Finished Goods 122750 =51250+71500
      Work in Process 122750